OpenStudy (anonymous):

X+2y=-6 in slope intercept form

Nnesha (nnesha):

ok so slope intercept form equation look like this y = mx + b where m is slope ( always with x) and y intercept so solve for y

Nnesha (nnesha):

subtract x both side first and let me know what you get after that

OpenStudy (anonymous):

I got 2y=-6

Nnesha (nnesha):

who kidnapped the x :???? where is our x???? :'(

Nnesha (nnesha):

when you subtract x on right side that should be stay there

OpenStudy (anonymous):

So I should still have x?

Nnesha (nnesha):

yes you do like i said slope intercept form is y =mx + b where m is slope always with x

Nnesha (nnesha):

\[2y = -x -6\] now you have this

Nnesha (nnesha):

now divide both side by 2

OpenStudy (anonymous):

Y=-x-3

Nnesha (nnesha):

who kidnapped my 2 now ????????????????????

Nnesha (nnesha):